OpenStudy (fairy733189):

what is 2 and 4/3 equal to? 3√8 2√8 3√16 2√16

OpenStudy (mathmale):

something's not quite right here. Are you asking for the square root of 2 4/3? Please double check to ensure that you've copied the problem down correctly.

OpenStudy (fairy733189):

Yes I did copy the problem down correctly

OpenStudy (mathmale):

If we read the problem literally: "what is 2 and 4/3 equal to?" we'd just convert it to an improper fraction. That would be 2 + 4/3 = 6/3 + 4/3 = 10/3. If, on the other hand, we were told to find the sqrt of 2 + 4/3, or the sqrt of 10/3, the correct result would be \[\sqrt{10/3}=\frac{ \sqrt(10 }{ \sqrt{3} }\]

OpenStudy (texaschic101):

none of those answers

OpenStudy (mathmale):

... which does not even come close to agreeing with any of the four answer choices. texaschic101: Any suggestions?

OpenStudy (texaschic101):

dunno...question copied right ?

OpenStudy (fairy733189):

Yes

OpenStudy (mathmale):

Let's suppose the answer were 3√8. This could be reduced. Sqrt(8) = Sqrt(2)Sqrt(4) = 2 Sqrt(2). Thus, we get 3*2Sqrt(2). This isn't even close to the original expression. Could you possibly move on, planning to ask your teacher about this problem later, and doing some new problems?

OpenStudy (welshfella):

is it possibly 

OpenStudy (welshfella):

this would simplify to